A train traveling at a speed of 30 miles per hour passes point A on its way to point B. At the same time, on a parallel track, another train traveling at a speed of 70 speeds per hour passes point B on its way to point A. If point A and point B are 300 miles apart, how far from from point B will these trains meet?
(additional question) does this question use the the formulas D=RT

Answer by ankor@dixie-net.com(22740)

A train traveling at a speed of 30 miles per hour passes point A on its way to point B. At the same time, on a parallel track, another train traveling at a speed of 70 speeds per hour passes point B on its way to point A. If point A and point B are 300 miles apart, how far from from point B will these trains meet?
(additional question) does this question use the the formulas D=RT
:
Using this formula:
:
Let t = the time required for this to happen
:
Train A dist + train B dist = 300 mi
30t + 70t = 300
100t = 300
t = 300/100
t = 3 hrs
:
They meet in 3 hrs; 3hrs from point B: 3 * 70 = 210 mi from point B